Here is the Agri-Data, mention the name of dependent and independent variables.

 

Part A: Prepare the line of regression Y on X.                                                    

Part B: Calculate r2 (coefficient of determination). Interpreted the byx.


District_2 Rapeseed & Mustard Rapeseed & Mustard Area

production in Metric Tones

UTTAR KASHI 437 873

CHAMOLI 193 407

RUDRA PRAYA 154 325

TEHRI GARHW 466 848

DEHRADUN 139 226




1
Expert's answer
2021-07-16T14:54:30-0400

Data



Rapeseed & Mustard production in Metric Tones is dependent variable = Y

Rapeseed mustard area production is independent variable = X

Part A

By using Excel



y=a + bx

=INTERCEPT(A2:A6,B2:B6)

a = -2.28

=SLOPE(A2:A6,B2:B6)

b= 0.52

y=0.52-2.28x

Part B

=RSQ(A2:A6,B2:B6)

"R^2=0.981"

The relation between Y and X is very strong. 98% of the variance of the dependent variable being studied is explained by the variance of the independent variable.

Graphical method with trend line.